# Asymptotic theory of in-context learning by linear attention

### v1

@@ Abstract
1 Introduction
Since their introduction by Vaswani et al. in 2017 [1], Transformers have become a cornerstone of modern artificial intelligence (AI). Originally designed for sequence modeling tasks, such as language modeling and machine translation, Transformers achieve state-of-the art performance across many domains, even those that are not inherently sequential [2]. Most strikingly, they underpin the breakthroughs achieved by large language models such as BERT [3], LLaMA [4], and the GPT series [5, 6, 7, 8].
@@ 1 Introduction
Despite many recent studies on understanding ICL, important questions about how and when ICL emerges in large language models are still mostly open. Large language models are trained (or pretrained) with a next token prediction objective. How do the different algorithmic and hyperparameter choices that go into the pretraining procedure affect ICL performance? What algorithms do Transformers implement for ICL? How many pretraining examples are required for ICL to emerge? How many examples should be provided within the input for the model to be able to solve an in-context task? How diverse should the tasks in the training dataset be for in-context learning of truly new tasks not observed in the training dataset?
In this paper, we address these questions by investigating the ICL capabilities of a linear attention module for linear regression tasks. This model setting allows us to derive an asymptotically precise theory of the learning curve. In the remainder of this section, we first provide a comprehensive overview of related works on ICL. Following this, we summarize our main contributions.
1.1 Related Works
@@
The striking ICL abilities of Transformers were thrust to the fore by Brown et al. [7]’s work on GPT-3. Focusing on natural language processing (NLP) tasks, they showed that ICL performance dramatically improves with an increase in the number of model parameters, with an increase in the number of examples in the model’s context, and with the addition of a natural language task description. In subsequent work, Wei et al. [11] proposed that the emergence of ICL with increasing scale is an abrupt, unpredictable transition. This perspective has substantially influenced proposed accounts for the emergence of ICL [12]. However, Schaeffer et al. [13] have disputed the idea that the emergence of ICL is unpredictable; they suggest that appropriately-chosen measures of otherwise hidden progress [14] reveal that ICL gradually develops with scale.
@@
Though ICL in NLP is both impressive and useful, these natural data do not allow precise experimental control. Towards a fine-grained understanding of the conditions required for ICL, many recent works have explored ICL of parametrically-controllable synthetic tasks, notably linear regression and classification. These works have identified various features of pretraining data distributions that contribute to the emergence of ICL [15, 16, 17, 18, 19]. Closely related to our work is a study of ICL of linear regression by Raventós et al. [18]. Their work identified a task diversity threshold for the emergence of ICL, below which a pretrained Transformer behaves as a Bayesian estimator with prior determined by the limited set of pretraining tasks. Above this threshold, the model’s performance matches that of within-context ridge regression, corresponding to a Gaussian prior over all tasks, including those not seen during pretraining. This work underscores the roles of task diversity, regularization, model capacity, and data structure in the emergence of ICL; a motivating objective of our work is to provide a theoretical account of their results.
@@
Many theoretical studies of ICL have centered on the idea that Transformers learn a particular algorithm during pretraining, which is then flexibly deployed to solve in-context tasks. In broad strokes, papers from this program of research often consider a particular algorithm for solving an in-context task, prove that Transformers can approximately implement this algorithm, and then empirically compare the ICL performance of a pre-trained Transformer to the performance of that algorithm [20, 21, 22, 23, 24, 25, 26, 27]. A clear consensus on which algorithm underlies ICL of linear regression in full transformers has yet to emerge [20, 21, 22, 23, 24, 25, 26, 27]. Within this line of research, closest to our work are a series of papers that consider ICL of linear regression by simplified Transformers using linear, rather than softmax, attention modules [27, 26, 23, 28, 25, 29, 30]. Zhang et al. [27] study these models in the limit of infinite pretraining dataset size (i.e., the population risk limit), and show that their performance on in-context linear regression nearly matches that of the Bayes-optimal estimator for the ICL task. However, they found that linear Transformers are not robust to shifts in the within-context covariate distribution. Zhang et al. [26] then showed that any optimizer of the within-context risk for a linear Transformer solves the ICL task with an approximation to one step of gradient descent from a learnable initialization, and that the resulting estimator can saturate the Bayes error for tasks with a Gaussian prior and non-zero mean. As we will discuss in Section 2, our reduction of the linear attention module is inspired in part by these works. In very recent work, Duraisamy [30] has studied the fine-sample risk of in-context linear regression with a single step of gradient descent, without directly analyzing Transformers. Ahn et al. [23] and Wu et al. [28] investigated how linear Transformers adapt to limited pretraining data and context length, again showing that in certain cases nearly-optimal error is achievable. Like these studies, our work considers linear attention, but our analysis, with its asymptotically sharp predictions of the ICL performance, allows us to pinpoint when and how the transition from memorization to ICL of linear regression occurs. We thus in closing highlight work by Reddy [19] on in-context classification, who analyzed the transition to ICL using a phenomenological model.
1.2 Summary of contributions
@@
We now summarize the primary contributions of our paper, relative to the prior art reviewed above. Following the recent literature, we focus on a simplified model of a Transformer that captures its key architectural motif: the linear self-attention module [25, 26, 29, 28, 27, 23]. Linear attention includes the quadratic interaction between inputs that lies at the heart of softmax attention, but does not include the normalization steps or fully-connected layers. This simplification makes the model more amenable for theoretical analysis. Our main result is a sharp asymptotic analysis of ICL of linear regression by linear attention, resulting in a more precisely predictive theory than previous population risk analyses or finite-sample bounds [27, 26]. The main contributions of our paper are structured as follows:
@@
    We begin in §2 by developing a simplified parameterization of linear self-attention that allows pretraining on the ICL linear regression task to be performed using ridge regression.
@@
    Within this simplified model, it is easy to identify the non-trivial scaling limit in which performance should be analyzed (§3): as the token dimension tends to infinity, the number of pretraining examples should scale quadratically (with the token dimension), while the context length and pretraining task diversity should scale linearly. In this joint limit, we can compute sharp asymptotics for ICL performance using random matrix theory.
@@
    Our theoretical results reveal several interesting phenomena (§3). First, we observe double-descent in the model’s ICL generalization performance as a function of pretraining dataset size, reflecting our assumption that it is pretrained to interpolation. Second, we uncover a transition to in-context learning as the pretraining task diversity increases. Concretely, there is a threshold task diversity above which linear attention saturates the Bayes error for the ICL regression task. Below that threshold, the model tends to memorize the limited set of pretraining tasks, and its excess risk is substantial. This transition recapitulates the empirical findings of Raventós et al. [18] in full Transformer models.
@@
    In §4, we show through numerical experiments that the insights from our theory derived on a simplified model transfer to full Transformer models with softmax self-attention. In particular, the scaling of pretraining sample complexity and task diversity with token dimension required for successful ICL is consistent.
More broadly, the study of solvable models is crucial for enhancing our grasp of how machine learning algorithms learn and generalize across various tasks. Understanding the mechanistic underpinnings of ICL of well-controlled synthetic tasks is an important prerequisite to understanding how it emerges from pretraining on natural data [19].
2 Problem formulation
We start by describing the setting of our study.
2.1 ICL of linear regression
@@
where $\epsilon_{i}$ is a Gaussian noise with mean zero and variance $\rho$, and $w\in\mathbb{R}^{d}$ is referred to as a task vector. We note that the task vector $w$ is fixed within a context, but can change between different contexts. The model has to learn $w$ from the $\ell$ pairs presented within the context, and use it to predict $y_{\ell+1}$ from $x_{\ell+1}$.
2.2 Linear self-attention
@@
For appropriately sized key, query, and value matrices $K,Q,V$, the output of a linear-attention block [32, 33, 31] is given by
@@
The output $A$ is a matrix while our goal is to predict a scalar, $y_{\ell+1}$. Following the choice of positional encoding in eq. 4, we will take $A_{d+1,\ell+1}$, the element of $A$ corresponding to the 00 prompt, as the prediction for $y_{\ell+1}$:
@@
2.3 Pretraining data
The model is pretrained on $n$ sample sequences, where the $\mu$th sample is a collection of $\ell+1$ vector-scalar pairs $\{x_{i}^{\mu}\in\mathbb{R}^{d},y_{i}^{\mu}\in\mathbb{R}\}_{i=1}^{\ell+1}$ related by the approximate linear mapping in eq. 1: $y_{i}^{\mu}=\langle x_{i}^{\mu},w^{\mu}\rangle+\epsilon_{i}^{\mu}$. Here, $w^{\mu}$ denotes the task vector associated with the $\mu$th sample.
We make the following statistical assumptions:
@@ 2.3 Pretraining data
    $x_{i}^{\mu}$ are $d$-dimensional random vectors, sampled i.i.d. over both $i$ and $\mu$ from a Gaussian distribution $\mathcal{N}(0,I_{d}/d)$.
@@
    For $1\leq\mu\leq n$, the task vector $w^{\mu}$ associated with the $\mu$th sample context is uniformly sampled from a finite set with $k$ elements, denoted by $\{w_{1},\ldots,w_{k}\}$. The elements of this set are independently drawn once at the beginning of training from
    $w_{i}\sim_{\text{i.i.d.}}\mathsf{Unif}(\mathcal{S}^{d-1}(\sqrt{d})),$ (7)
    where $\mathsf{Unif}(\mathcal{S}^{d-1}(\sqrt{d}))$ denotes the uniform distribution on the sphere $\mathcal{S}^{d-1}(\sqrt{d})$ of radius $\sqrt{d}$. The variable $k$ controls the task diversity in the pretraining data set. Importantly, $k$ can be less than $n$, in which case the same task vector may be repeated multiple times.
@@
2.4 Parameter reduction
Before specifying a training procedure, it is insightful to examine the prediction mechanism of the linear attention module for the ICL task. This turns out to be a fruitful exercise, shedding light on critical questions: Can linear self-attention learn linear regression in-context? If so, what information do model parameters learn from data in solving this ICL problem? By closely examining these aspects, we can also formulate a simplified problem that lends itself to analytical study.
We start by rewriting the output of the linear attention module, eq. 6, in an alternative form. Following Zhang et al. [27], we define
@@ 2.4 Parameter reduction
where $V_{11}\in\mathbb{R}^{d\times d}$, $v_{12},v_{21}\in\mathbb{R}^{d\times 1}$, $v_{22}\in\mathbb{R}$, $M_{11}\in\mathbb{R}^{d\times d}$, $m_{12},m_{21}\in\mathbb{R}^{d\times 1}$, and $m_{22}\in\mathbb{R}$. The expression we desire is
$\hat{y}$ $=\frac{1}{\ell}\left\langle{x_{\ell+1},v_{22}M_{11}^{\top}\sum_{i=1}^{\ell}y_{i}x_{i}+v_{22}m_{21}\sum_{i=1}^{\ell}y_{i}^{2}+M_{11}^{\top}\sum_{i=1}^{\ell+1}x_{i}x_{i}^{\top}v_{21}+m_{21}\sum_{i=1}^{\ell}y_{i}x_{i}^{\top}v_{21}}\right\rangle,$ (13)
@@
This expression reveals several interesting points. First, not all parameters in (12) contribute to the output: We can ignore all the parameters except the last row of $V$ and the first $d$ columns of $M$. Second, the first term
@@
offers a hint about how the linear attention module might be solving the task. The sum $\frac{1}{\ell}\sum_{i\leq\ell}y_{i}x_{i}$ is a noisy estimate of $\mathbb{E}[xx^{\top}]w$ for that context. Hence, if the parameters of the model are such that $v_{22}M_{11}^{\top}$ is approximately $\mathbb{E}[xx^{\top}]^{-1}$, this term alone makes a good prediction for the output. Third, the third term does not depend on outputs $y$, and thus does not directly contribute to the ICL task that relies on the relationship between $x$ and $y$. Fourth, the last term only considers a one dimensional projection of $x$ onto $v_{21}$. Because the task vectors $w$ and $x$ are isotropic in the statistical models that we consider, there are no special directions in the problem. Consequently, we expect the optimal $v_{21}$ to be approximately zero by symmetry considerations.
@@
We note that Zhang et al. [27] provide an analysis of population risk (whereas we focus on empirical risk) for a related reduced model in which they set $v_{21}=0$ and $m_{21}=0$. Consequently, the predictors they study differ from ours (17) by an additive term. They justify this choice through an optimization argument: if these parameters are initialized to zero, they remain zero under gradient descent optimization of the population risk, given certain conditions.

### v2

@@ Abstract
I Introduction
Since their introduction by Vaswani et al. in 2017 [1], Transformers have become a cornerstone of modern artificial intelligence (AI). Originally designed for sequence modeling tasks, such as language modeling and machine translation, Transformers achieve state-of-the art performance across many domains, even those that are not inherently sequential [2]. Most strikingly, they underpin the breakthroughs achieved by large language models such as BERT [3], LLaMA [4], the GPT series [5, 6, 7, 8], the Claude model family [9], and DeepSeek R1 [10].
@@ I Introduction
Despite recent progress in understanding ICL, fundamental questions about when and how ICL emerges in large language models remain unresolved. These models are typically trained or pretrained using a next-token prediction objective, but the impact of various algorithmic and hyperparameter choices during pretraining on ICL performance is still not well understood. What mechanisms do Transformers implement to facilitate ICL? How many pretraining examples are necessary for ICL capabilities to arise? Furthermore, how many examples must be provided within the input context for the model to successfully perform an in-context task? Another important question is the degree of task diversity in the training data: how diverse must the training tasks be to enable ICL for truly novel tasks that the model has not encountered before?
In this paper, we address these questions by investigating the ICL capabilities of a linear attention module for linear regression tasks. This simplified model setting, which exhibits the minimal architectural features to perform ICL for linear regression, enables us to derive an asymptotically precise theory of ICL performance, elucidating its exact dependence on various hyperparameters. In the remainder of this section, we first provide an overview of related work on ICL. We then summarize our main contributions.
I.1 Related Work
@@
The striking ICL abilities of Transformers were thrust to the fore by Brown et al. [7]’s work on GPT-3. Focusing on natural language processing (NLP) tasks, they showed that ICL performance dramatically improves with an increase in the number of model parameters, with an increase in the number of examples in the model’s context, and with the addition of a natural language task description. In subsequent work, Wei et al. [13] proposed that the emergence of ICL with increasing scale is an abrupt, unpredictable transition. This perspective has substantially influenced proposed theories for the emergence of ICL [14]. However, Schaeffer et al. [15] have challenged the idea that the emergence of ICL is unpredictable; they suggest that appropriately-chosen measures of otherwise hidden progress [16] reveal that ICL gradually develops with scale.
@@
Though ICL in NLP is both impressive and useful, these natural data do not allow precise experimental control and study. Towards a fine-grained understanding of the conditions required for ICL, many recent works have explored ICL of parametrically-controllable synthetic tasks, notably linear regression and classification. These works have identified various features of pretraining data distributions that contribute to the emergence of ICL [17, 18, 19, 20, 21]. Closely related to our work is a study of ICL of linear regression by Raventós et al. [20]. Their work identified a task diversity threshold for the emergence of ICL, below which a pretrained Transformer behaves as a Bayesian estimator with a prior determined by the limited set of pretraining tasks. Above this threshold, the model’s performance approaches that of within-context optimal Bayesian ridge regression, corresponding to a Gaussian prior over all tasks, including those not seen during pretraining. A motivating objective of our work is to provide a theoretical account of the empirical findings made by Raventós et al. [20], which underscore the roles of task diversity, regularization, model capacity, and data structure in the emergence of ICL.
@@
Many theoretical studies of ICL have centered on the idea that Transformers learn a particular algorithm during pretraining, which is then flexibly deployed to solve in-context tasks. In broad strokes, papers from this program of research often consider a particular algorithm for solving an in-context task, prove that Transformers can approximately implement this algorithm, and then empirically compare the ICL performance of a pre-trained Transformer to the performance of that algorithm [22, 23, 24, 25, 26, 27, 28, 29]. A clear consensus on which algorithm underlies ICL of linear regression in full transformers has yet to emerge [22, 23, 24, 25, 26, 27, 28, 29]. Within this line of research, closest to our work are a series of papers that consider ICL of linear regression by simplified Transformers using linear, rather than softmax, attention modules [29, 28, 25, 30, 27, 31, 32]. Zhang, Frei, and Bartlett [29] studied these models in the limit of infinite pretraining dataset size (i.e., the population risk limit), and show that their performance on in-context linear regression nearly matches that of the Bayes-optimal estimator for the ICL task. However, they found that linear Transformers are not robust to shifts in the within-context covariate distribution. Zhang, Wu, and Bartlett [28] then showed that any optimizer of the within-context risk for a linear Transformer solves the ICL task with an approximation to one step of gradient descent from a learnable initialization, and that the resulting estimator can saturate the Bayes error for tasks with a Gaussian prior and non-zero mean. As we will discuss in Section II, our reduction of the linear attention module is inspired in part by these works. In very recent work, Duraisamy [32] has studied the fine-sample risk of in-context linear regression with a single step of gradient descent, without directly analyzing Transformers. An et al. [25] and Wu et al. [30] investigated how linear Transformers adapt to limited pretraining data and context length, again showing that in certain cases nearly-optimal error is achievable. Like these studies, our work considers linear attention; but our analysis, with its asymptotically sharp predictions, allows us to elucidate the exact dependence of ICL performance on various hyperparameters and to pinpoint when and how the transition from memorization to ICL of linear regression occurs. In closing, we highlight the work of Reddy [21], who analyzed the emergence of in-context classification through a phenomenological model.
I.2 Summary of contributions
@@
We now summarize the main contributions of our paper relative to the prior art reviewed above. Building on recent literature, we focus on a simplified model of a Transformer that captures its key architectural motif: the linear self-attention module [27, 28, 31, 30, 29, 25]. Linear attention includes the quadratic pairwise interactions between inputs that lie at the heart of softmax attention, but it omits the normalization steps and fully connected layers. This simplification makes the model more amenable to theoretical analysis. Our main result is a sharp asymptotic analysis of ICL for linear regression using linear attention, leading to a more precisely predictive theory than previous population risk analyses or finite-sample bounds [29, 28]. The main contributions of our paper are structured as follows:
@@
    We begin in §II by developing a simplified parameterization of linear self-attention that allows pretraining on the ICL linear regression task to be performed using ridge regression.
@@
    Within this simplified model, we identify a phenomenologically rich scaling limit in which the ICL performance can be analyzed (§III). As the token dimension tends to infinity, we allow the number of pretraining examples scale quadratically with the token dimension, while the context length and pretraining task diversity scale linearly. In this joint limit, we compute sharp asymptotics for ICL performance using random matrix theory.
@@
    The asymptotically precise theory curves we derive reveal several interesting phenomena (§III). First, we observe double-descent in the model’s ICL generalization performance as a function of pretraining dataset size, reflecting our assumption that it is pretrained to interpolation. Second, we study the non-monotonic dependence of ICL performance on context length. Last, we uncover a transition from memorization to in-context learning as the pretraining task diversity increases. This transition recapitulates the empirical findings of [20] in full Transformer models.
@@
    In §IV, we demonstrate through numerical experiments that the insights from our theory, derived using the simplified linear attention model, transfer to full Transformer models with softmax self-attention. In particular, the scaling of pretraining sample complexity and task diversity with token dimension required for successful ICL is consistent.
II Problem formulation
We begin by describing the setting of our study.
II.1 ICL of linear regression
@@
where $\epsilon_{i}$ is a Gaussian noise $w\in\mathbb{R}^{d}$ is referred to as a task vector. We note that the task vector $w$ is fixed within a context, but can change between different contexts. The model has to learn $w$ from the $\ell$ pairs presented within the context, and use it to predict $y_{\ell+1}$ from $x_{\ell+1}$.
II.2 Linear self-attention
@@
For value matrix $V\in\mathbb{R}^{(d+1)\times(d+1)}$ and key and query matrices $K,Q$ such that $K^{\top}Q\in\mathbb{R}^{(d+1)\times(d+1)}$, the output of a linear-attention block [34, 35, 33] is given by
@@
The output $A$ is a matrix while our goal is to predict a scalar, $y_{\ell+1}$. Following the choice of positional encoding in (4), we will take $A_{d+1,\ell+1}$, the element of $A$ corresponding to the 00 prompt, as the prediction for $y_{\ell+1}$:
@@
II.3 Pretraining data
The model is pretrained on $n$ sample sequences, where the $\mu$th sample is a collection of $\ell+1$ vector-scalar pairs $\{x_{i}^{\mu}\in\mathbb{R}^{d},y_{i}^{\mu}\in\mathbb{R}\}_{i=1}^{\ell+1}$ related by the approximate linear mapping in (1): $y_{i}^{\mu}=\langle x_{i}^{\mu},w^{\mu}\rangle+\epsilon_{i}^{\mu}$. Here, $w^{\mu}$ denotes the task vector associated with the $\mu$th sample. We make the following statistical assumptions:
@@ II.3 Pretraining data
    $x_{i}^{\mu}$ are $d$-dimensional random vectors, sampled i.i.d. over both $i$ and $\mu$ from an isotropic Gaussian distribution $\mathcal{N}(0,I_{d}/d)$.
@@
    At the beginning of training, construct a finite set with $k$ task vectors, denoted by
    $\Omega_{k}=\left\{w_{1},w_{2},\ldots,w_{k}\right\}.$ (7)
    The elements of this set are independently drawn once from
    $w_{i}\sim_{\text{i.i.d.}}\mathcal{N}(0,I_{d}).$ (8)
    For $1\leq\mu\leq n$, the task vector $w^{\mu}$ associated with the $\mu$th sample context is uniformly sampled from $\Omega_{k}$. Note that the variable $k$ controls the task diversity in the pretraining dataset. Importantly, $k$ can be less than $n$, in which case the same task vector from $\Omega_{k}$ will be repeated multiple times.
@@
II.4 Parameter reduction
Before specifying the training procedure, it is insightful to first examine the prediction mechanism of the linear attention module for the ICL task. This proves to be a fruitful exercise, shedding light on key questions: Can linear self-attention learn linear regression in-context? If so, what do the model parameters learn from the data to solve this ICL problem? By closely analyzing these aspects, we can also formulate a simplified problem that lends itself to analytical study.
We start by rewriting the output of the linear attention module, (6), in an alternative form. Following [29], we define
@@ II.4 Parameter reduction
where $V_{11}\in\mathbb{R}^{d\times d}$, $v_{12},v_{21}\in\mathbb{R}^{d}$, $v_{22}\in\mathbb{R}$, $M_{11}\in\mathbb{R}^{d\times d}$, $m_{12},m_{21}\in\mathbb{R}^{d}$, and $m_{22}\in\mathbb{R}$. We assume that the inner dimension of $K^{\top}Q$ is greater than equal to $d+1$ so that matrix $M$ can achieve full rank. From (5) and (6), one can check that
$\hat{y}=\frac{1}{\ell}\big{\langle}x_{\ell+1},v_{22}M_{11}^{\top}\sum_{i=1}^{\ell}y_{i}x_{i}+v_{22}m_{21}\sum_{i=1}^{\ell}y_{i}^{2}+M_{11}^{\top}\sum_{i=1}^{\ell+1}x_{i}x_{i}^{\top}v_{21}+m_{21}\sum_{i=1}^{\ell}y_{i}x_{i}^{\top}v_{21}\big{\rangle},$ (14)
@@
This expression reveals several interesting points, including how this model could express a solution to the ICL task. First, not all parameters in (13) contribute to the output: we can discard all the parameters except for the last row of $V$ and the first $d$ columns of $M$. Second, the first term
@@
offers a hint about how the linear attention module might be solving the task. The sum $\frac{1}{\ell}\sum_{i\leq\ell}y_{i}x_{i}$ is a noisy estimate of $\mathbb{E}[xx^{\top}]w$ for that context. Hence, if the parameters of the model are such that $v_{22}M_{11}^{\top}$ is approximately $\mathbb{E}[xx^{\top}]^{-1}$, this term alone makes a good prediction for the output. Third, the third term, $M_{11}^{\top}\sum_{i=1}^{\ell+1}x_{i}x_{i}^{\top}v_{21}$ does not depend on outputs $y$, and thus does not directly contribute to the ICL task that relies on the relationship between $x$ and $y$. Finally, the fourth term, $m_{21}\sum_{i=1}^{\ell}y_{i}x_{i}^{\top}v_{21}$, only considers a one dimensional projection of $x$ onto $v_{21}$. Because the task vectors $w$ and $x$ are isotropic in the statistical models that we consider, there are no special directions in the problem. Consequently, we expect the optimal $v_{21}$ to be approximately zero by symmetry considerations.
@@
We note that [29] provides an analysis of the population risk (whereas we focus on empirical risk) for a related reduced model in which both $v_{21}$ and $m_{21}$ are set to 00. Consequently, the predictors considered in [29] slightly differ from ours in (16)–(18) by an additive term (due to $m_{21}$). The authors of [29] justify this reduced model through an optimization argument: if these parameters $v_{21}$ and $m_{21}$ are initialized to zero, they remain zero under gradient descent optimization of the population risk.